OpenStudy (anonymous):

perform the indicated operations then simplify each radical... 13 sqrt 2 - 5 sqrt 8 + 2 sqrt32 help!!

OpenStudy (anonymous):

\[13\sqrt{2}-5 \sqrt{8}+2 \sqrt{32}\]

OpenStudy (anonymous):

5sqrt8 = 10sqrt2 2sqrt32 = 8sqrt2 13sqrt2 - 10sqrt2 - 8sqrt2 = -5sqrt2

OpenStudy (anonymous):

see.. sqrt 8 = sqrt (2 x 2 x 2) = sqrt(4) x sqrt(2) = 2sqrt(2)

OpenStudy (anonymous):

@ynapot97 ?